# Analytical results for the Quantum Alternating Operator Ansatz with Grover Mixer

### v2

@@ Abstract
The Grover mixer operator is a variational version of Grover’s diffusion operator, introduced as a mixing operator for the Quantum Alternating Operator Ansatz (QAOA) and used in a variant known as Grover Mixer QAOA (GM-QAOA). An important property of QAOA with Grover mixer is that its expectation value is invariant over any permutation of states. As a consequence, the algorithm is independent of the structure of the problem. If, on the one hand, this characteristic raises serious doubts about the capacity of the algorithm to overcome the bound of the unstructured search problem, on the other hand, it can pave the way to its analytical study. In this sense, a prior work introduced a statistical approach to analyze GM-QAOA that results in an analytical expression for the expectation value depending on the probability distribution associated with the problem Hamiltonian spectrum. Although the method provides surprising simplifications in calculations, the expression depends exponentially on the number of layers, which makes direct analytical treatment unfeasible. In this work, we extend the analysis to the more simple context of Grover Mixer Threshold QAOA (GM-Th-QAOA), a variant that replaces the phase separation operator of GM-QAOA to encode a threshold function. As a result, we obtain an expression for the expectation value independent of the number of layers and, with it, we provide bounds for different performance metrics. Furthermore, we extend the analysis to a more general context of QAOA with Grover mixer, which we called Grover-based QAOA. In that framework, which allows the phase separation operator to encode any compilation of the cost function, we generalize all the bounds by using a contradiction argument with the optimality of Grover’s algorithm on the unstructured search problem. As a result, we get the main contribution of this work, an asymptotic lower bound on the quantile achieved by the expectation value that formalizes the notion that the Grover mixer, at most, reflects a quadratic Grover-style speed-up over classical brute force. We apply that bound on the Max-Cut problem to the particular class of complete bipartite graphs and argue that the number of rounds required to achieve a fixed approximation ratio must grow exponentially with the number of vertices/edges, a severe limitation on the performance of the algorithm.
@@ I Introduction
The original mixing operator of QAOA uses the transverse field mixer Hamiltonian [4], which is given by a sum of Pauli-$X$ operators. Since then, many other variations with different types of mixers have already been introduced in the literature [5, 8, 9, 10, 11]. There is numerical evidence that the choice of mixing operator significantly affects the performance of QAOA [12, 13, 14] and therefore choosing the ideal mixer for a given optimization problem is an important research topic. One variant of particular interest is the Grover Mixer Quantum Alternating Operator Ansatz (GM-QAOA), introduced for both unconstrained [12, 15, 16] and constrained [8] optimization context. The mixing operator of GM-QAOA is a variational version of Grover’s diffusion operator [17, 18] called Grover mixer operator. In GM-QAOA formulation, a necessary condition to the construction of the mixing operator is the existence of an efficient preparation of uniform superposition over the feasible states—which covers constrained problems like the Traveling Salesman Problem, the $k$-Vertex-Cover, and the Discrete Portfolio Rebalancing. Alternatively, the operator of Grover mixer can be constructed with the formulation of Quantum Walk-based Optimization Algorithm (QWOA) [9, 10], a generalization of QAOA that interprets the mixing operator as a continuous-time quantum walk operator [19, 20, 21]. In that case, the Grover mixer operator is equivalent to QWOA on the complete graph up to a change of scale on the operator parameter. Problems such as the Capacitated Vehicle Routing [22] and the Portfolio Optimization [23] have already been numerically studied within the QWOA framework.
@@
Another variant of Quantum Alternating Operator Ansatz using the Grover mixer is the so-called Grover Mixer Threshold QAOA (GM-Th-QAOA) [24], an algorithm combing such mixing operator with the more general Threshold QAOA, which in turn changes the original phase separation to encode a compilation of the cost function into a threshold function splitting the solution space from a value. In particular, the choice of all angles as being equal to $\pi$ reduces the GM-Th-QAOA to a Grover’s search for marked states above (considering the original definition) a given threshold. An advantage of this variant is admitting an efficient procedure to find optimal parameters—the angles and the threshold value—that eliminates the costly outer loop parameter finding of the usual QAOA. Furthermore, it has been numerically observed that the performance of GM-Th-QAOA consistently overcomes GM-QAOA in all instances considered [24, 13, 14].
@@
The performance of QAOA with the Grover mixer, individually or compared with other mixers, has already been considered in the literature. The initial thought, corroborated by numerical experiments on small instances, was that the Grover mixer would overcome transverse field mixer due to its ability to mix quickly and its global symmetry among states [12, 10, 23]. However, later experiments on larger instances indicated that this advantage soon disappeared with Grover mixer losing to transverse field mixer on unconstrained context [13] and performing even exponentially worse than the clique mixer [25, 26] on constrained problems [14]. One can argue that the worst performance of the Grover mixer may be due to the fact that it depends only on the distribution of the solution space—by the global symmetry, the Grover mixer is invariant under any permutation of states [27]—so the algorithm does not see the structure of the optimization problem and possibly is limited to the bound of the unstructured search problem [28] (i.e., a Grover-style quadratic speed-up), drastically compromising algorithm performance on large instances. On the other hand, other mixers, such as transverse field and clique, could, in principle, overcome that limit by exploiting the underlying problem structure.
@@
Although the Headley and Wilhelm [27] method provides surprising simplifications in GM-QAOA expectation value calculations, the expression is still too complicated to obtain formal bounds on the algorithm’s performance through direct analytical treatment. In the present work, with the motivation of understanding the theoretical potential of the Grover mixer on combinatorial optimization—especially to investigate the issue concerning the quadratic Grover-like speed-up—we extended the analysis to the more simple case of GM-Th-QAOA. Using the well-known formula of probability of Grover’s algorithm and its optimality on average probability for the unstructured search [35, 36], we provide an expression for the expectation value with complexity independent on the number of layers, which allows to study the asymptotic behavior of the algorithm. Rather than the characteristic function, the prominent statistical quantity here is the conditional expected value. With a closed-form expression, we prove the conjecture on which the efficient method of Golden et al. [24] of finding the optimization parameters is based (see details on Subsec. III.1). Furthermore, we provide bounds on the performance of the expectation value of GM-Th-QAOA with the statistical quantities of quantile and the standard score. On the first, we get an asymptotic tight bound that implies the expected quadratic Grover-style speed-up of GM-Th-QAOA over classical brute force. On the second, we conclude that the maximum standard score achieved by the expectation value of GM-Th-QAOA, hit by binary functions with specific ratios, scales linearly with the number of rounds. As an immediate consequence, we bound the number of layers to achieve a fixed approximation ratio. Finally, we combine both bounds to argue that the algorithm’s performance is closely related to the asymptotic behavior of the probability distribution on the limit of its support.
@@
To get stronger results about the Grover mixer, we also consider a more general version of QAOA with the Grover mixer that we called Grover-based QAOA. In that framework, which includes both GM-QAOA and GM-Th-QAOA, is allowed that the phase separation operator codifies any compilation of the cost function in a real-valued function. We generalize all bounds of GM-Th-QAOA to Grover-based QAOA with a technique that consists of bounding the maximum amplification of the probability of measuring a set of degenerate states. The argument used for this is based on showing that if there is an amplification greater than the one provided by Grover’s search, we contradict the optimality of average provability on unstructured search [35, 36], building an explicit algorithm that performance better than Grover’s algorithm. All bounds on Grover-based QAOA follow the same asymptotic behavior as its correspondents on GM-Th-QAOA. As a consequence, we get the principal contribution of this work, the formalization of the notion that the Grover mixer is limited to the quadratic speed-up over classical brute force with an asymptotic performance, for instance, analog to the Grover Adaptive Search (GAS) [37, 38, 39, 40]. We apply the bounds in the context of the Max-Cut problem. That way, by using the knowledge of the asymptotic behavior of the probability distribution associated with the particular case of complete bipartite graphs, we argue that for this class of graphs, the number of rounds required to achieve a fixed approximation ratio must grow exponentially with the number of vertices/edges, a severe limitation on the performance of the Grover mixer. More than that, the construction suggests that it is likely that the same happens with other classes of graphs and even with other combinatorial optimization problems.
@@ II.1 The Grover-based QAOA
for a threshold value $t$ that must be optimized. The minus sign on Eq. (2) is because in this work we consider QAOA for minimization problems. The expectation value of GM-QAOA and GM-Th-QAOA of depth $r$ is denoted here by $E_{r}(\bm{\beta},\bm{\gamma})$ and $E_{r}(t)$, respectively, while $E_{r}$ denotes the expectation value of a generic Grover-based QAOA.
@@ II.2 Random variables
Let $X$ be the random variable of uniformly sampling an element on the set $S$ and calculating the cost function. The function $f_{X}(x)=|\{k\in S:c(k)=x\}|/M$ is the probability mass function (pmf) of $X$, and the support $R_{X}$ of $X$ is a countable subset of real numbers. We denote the mean, standard deviation, minimum value, and maximum value of $X$ by $\mu=\operatorname{E}[X]$, $\sigma=\sqrt{\operatorname{E}[X-\mu]^{2}}$, $R_{X}^{min}$, and $R_{X}^{max}$, respectively, and assume $0<\sigma<\infty$ (ignoring the degenerate distribution). Provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$, the approximation ratio is defined by $\lambda=E_{r}/R_{X}^{min}$. We also define the random variable $Y$ as $Y=X-\mu$, and the standard random variable $Z$ as $Z=(X-\mu)/\sigma$.
The main statistical quantity of the analysis of GM-Th-QAOA is the conditional expectation. The expectation of $X$ given $X\leq x$ and the expectation of $X$ given $X>x$ are
@@
respectively, where $G_{X}(x)=\sum_{k\in R_{X}:k\leq x}kf_{X}(k)$ and $F_{X}(x)=\sum_{k\in R_{X}:k\leq x}f_{X}(k)$ is the cumulative distribution function (cdf). Note that with these definitions, we can consider any $x\in\mathbb{R}$ as argument.
Furthermore, in part of the present work, $F_{X}(x)$ and $G_{X}(x)$ must be differentiable. However, $X$ was defined to be discrete. To bypass that obstacle, note that we can write $F_{X}(x)$ and $G_{X}(x)$ as
$\begin{split}&F_{X}(x)=\sum_{k\in R_{X}}f_{X}(k)\theta(x-k),\\ &G_{X}(x)=\sum_{k\in R_{X}}kf_{X}(k)\theta(x-k),\end{split}$ (4)
where $\theta(x)$ is the Heaviside step function. The derivatives of $F_{X}(x)$ and $G_{X}(x)$ are then given by $f_{X}^{G}(x)$ and $xf_{X}^{G}(x)$ respectively, where
$\begin{split}f_{X}^{G}(x)=\sum_{k\in R_{X}}f_{X}(k)\delta(x-k),\end{split}$ (5)
with $\delta(x)$ as the Dirac delta function. The distribution $f_{X}^{G}(x)$ is an extension of the concept of probability density function for non-continuous random variables, usually called generalized probability density function [41].
In other situations, on the other hand, it is convenient for $F_{X}(x)$ to be continuous, which does not occur for generalized density functions. That way, we consider $X$ as a continuous random variable with an asymptotic approximation, in the same way as Headley and Wilhelm [27] approach. Consequently, all summations are replaced by integrals, and the probability mass function becomes a probability density function (pdf). However, beyond mere convenience, this assumption is justifiable since the main QAOA target is NP-Hard optimization problems with sufficient scalability to make it such a good approximation.

### v3

@@ Abstract
The Grover mixer operator is a variational version of Grover’s diffusion operator, introduced as a mixing operator for the Quantum Alternating Operator Ansatz (QAOA) and used in a variant known as Grover Mixer QAOA (GM-QAOA). An important property of QAOA with Grover mixer is that its expectation value is invariant over any permutation of states. As a consequence, the algorithm is independent of the structure of the problem. If, on the one hand, this characteristic raises serious doubts about the capacity of the algorithm to overcome the bound of the unstructured search problem, on the other hand, it can pave the way to its analytical study. In this sense, a prior work introduced a statistical approach to analyze GM-QAOA that results in an analytical expression for the expectation value depending on the probability distribution associated with the problem Hamiltonian spectrum. Although the method provides surprising simplifications in calculations, the expression depends exponentially on the number of layers, which makes direct analytical treatment unfeasible. In this work, we extend the analysis to the more simple context of Grover Mixer Threshold QAOA (GM-Th-QAOA), a variant that replaces the phase separation operator of GM-QAOA to encode a threshold function. As a result, we obtain an expression for the expectation value independent of the number of layers and, with it, we provide bounds for different performance metrics. Furthermore, we extend the analysis to a more general context of QAOA with Grover mixer, which we called Grover-based QAOA. In that framework, which allows the phase separation operator to encode any compilation of the cost function, we generalize all the bounds by using an argument by contradiction with the optimality of Grover’s algorithm on the unstructured search problem. As a result, we get the main contribution of this work, an asymptotic lower bound on the quantile achieved by the expectation value that formalizes the notion that the Grover mixer, at most, reflects a quadratic Grover-style speed-up over classical brute force. We apply that bound on the Max-Cut problem to the particular class of complete bipartite graphs and argue that the number of rounds required to achieve guarantees for any approximation ratio must grow exponentially with the number of vertices/edges, a severe limitation on the performance of the algorithm.
@@ I Introduction
The original mixing operator of QAOA uses the transverse field mixer Hamiltonian [4], which is given by a sum of Pauli-$X$ operators. Since then, many other variations with different types of mixers have already been introduced in the literature [5, 8, 9, 10, 11]. There is numerical evidence that the choice of mixing operator significantly affects the performance of QAOA [12, 13, 14] and therefore choosing the ideal mixer for a given optimization problem is an important research topic. One variant of particular interest is the Grover Mixer Quantum Alternating Operator Ansatz (GM-QAOA), introduced for both unconstrained [12, 15, 16] and constrained [8] optimization context. The mixing operator of GM-QAOA is a variational version of Grover’s diffusion operator [17, 18] called Grover mixer operator. In GM-QAOA formulation, a necessary condition to the construction of the mixing operator is the existence of an efficient preparation of uniform superposition over the feasible states—which covers constrained problems like the Traveling Salesman Problem, the Max $k$-Vertex Cover, and the Discrete Portfolio Rebalancing. Alternatively, the operator of Grover mixer can be constructed with the formulation of Quantum Walk-based Optimization Algorithm (QWOA) [9, 10], a generalization of QAOA that interprets the mixing operator as a continuous-time quantum walk operator [19, 20, 21]. In that case, the Grover mixer operator is equivalent to QWOA on the complete graph up to a change of scale on the operator parameter. Problems such as the Capacitated Vehicle Routing [22] and the Portfolio Optimization [23] have already been numerically studied within the QWOA framework.
@@
Another variant of Quantum Alternating Operator Ansatz using the Grover mixer is the so-called Grover Mixer Threshold QAOA (GM-Th-QAOA) [24], an algorithm combing such mixing operator with the more general Threshold QAOA, which in turn changes the original phase separation to encode a compilation of the cost function into a threshold function splitting the solution space from a value. That variant is closely related to Grover’s algorithm. In the general context of QAOA, Jiang, Rieffel, and Wang [25] show that the algorithm with original transverse field mixer can achieve the same asymptotic quadratic speed-up over classical brute force as the Grover’s algorithm for the unstructured search problem. In the particular case of GM-Th-QAOA, the relationship is much more direct since the choice of all angles as being equal to $\pi$ reduces the algorithm to a Grover’s search for marked states above (considering the original definition) a given threshold. An advantage of this variant is admitting an efficient procedure to find optimal parameters—the angles and the threshold value—that eliminates the costly outer loop parameter finding of the usual QAOA. Furthermore, it has been numerically observed that the performance of GM-Th-QAOA consistently overcomes GM-QAOA in all instances considered [24, 13, 14].
@@
The performance of QAOA with the Grover mixer, individually or compared with other mixers, has already been considered in the literature. The initial thought, corroborated by numerical experiments on small instances, was that the Grover mixer would overcome transverse field mixer due to its ability to mix quickly and its global symmetry among states [12, 10, 23]. However, later experiments on larger instances indicated that this advantage soon disappeared with Grover mixer losing to transverse field mixer on unconstrained context [13] and performing even exponentially worse than the clique mixer [26, 27] on constrained problems [14]. One can argue that the worst performance of the Grover mixer may be due to the fact that it depends only on the distribution of the solution space—by the global symmetry, the Grover mixer is invariant under any permutation of states [28]—so the algorithm does not see the structure of the optimization problem and is possibly limited to the bound of the unstructured search problem [29] (i.e., a Grover-style quadratic speed-up), drastically compromising algorithm performance on large instances. On the other hand, other mixers, such as transverse field and clique, could, in principle, overcome that limit by exploiting the underlying problem structure.
@@ I.1 Our contributions
Although the Headley and Wilhelm [28] method provides surprising simplifications in GM-QAOA expectation value calculations, the expression is still too complicated to obtain formal bounds on the algorithm’s performance through direct analytical treatment. In the present work, with the motivation of understanding the theoretical potential of the Grover mixer on combinatorial optimization—mainly to investigate the issue concerning the quadratic Grover-like speed-up—we extended the analysis to the more simple case of GM-Th-QAOA. Using the well-known formula of the probability of Grover’s algorithm and its optimality on average probability for the unstructured search [36, 37], we provide an expression for the expectation value in Theorem 1 with complexity independent of the number of layers, which allows us to study the asymptotic behavior of the algorithm. Rather than the characteristic function, the prominent statistical quantity here is the conditional expected value. With a closed-form expression, in Theorem 2, we prove the conjecture on which the efficient method of Golden et al. [24] of finding the optimization parameters is based (see details on Subsec. III.1). Furthermore, we provide bounds on the performance of the expectation value of GM-Th-QAOA with the statistical quantities of quantile and the standard score. On the first, we get in Theorem 3 an asymptotic tight bound that implies the expected quadratic Grover-style speed-up of GM-Th-QAOA over classical brute force, the most relevant contribution of this work to the specific context of GM-Th-QAOA. On the second, from Lemma 1, we conclude that the maximum standard score achieved by the expectation value of GM-Th-QAOA is hit by binary functions with specific ratios, and investigating the performance of these distributions, we state in Theorem 4 that the standard score scales at most linearly with the number of rounds. As an immediate consequence, we bound in Corollary 3 the minimum number of layers to achieve guarantees for a fixed approximation ratio. Finally, we combine both bounds to argue that the algorithm’s performance is closely related to the asymptotic behavior of the probability distribution on the limit of its support.
@@
To get stronger results about the Grover mixer, we also consider a more general version of QAOA with the Grover mixer that we called Grover-based QAOA. In that framework, which includes both GM-QAOA and GM-Th-QAOA, is allowed that the phase separation operator codifies any compilation of the cost function in a real-valued function. We generalize all bounds of GM-Th-QAOA to Grover-based QAOA with a technique used in Lemma 2 that consists of bounding the maximum amplification of the probability of measuring a set of degenerate states. The argument used for this consists of showing that if there is an amplification of states greater than the one provided by Grover’s search, i.e., a quadratic amplification on the probability of sampled states, we contradict the optimality of average provability on unstructured search [36, 37], building an explicit algorithm that performance better than Grover’s algorithm. With that limitation on the probability amplification of states, we get in Theorem 5 a bound on the expectation value of any Grover-based QAOA, and from it, we conclude that all bounds on Grover-based QAOA follow the same asymptotic behavior as its correspondents on GM-Th-QAOA. Specifically, Corollary 4 generalize Theorem 3 and Theorem 6 is the analogous for Grover-based QAOA of both Theorem 4 and Corollary 3. In particular, Corollary 4 gives the main general contribution of this work, the formalization of the notion that the Grover mixer is limited to the quadratic speed-up over classical brute force with an asymptotic performance, for instance, analog to the Grover Adaptive Search (GAS) [38, 39, 40, 41]. We apply the bounds in the context of the Max-Cut problem. That way, by using the knowledge of the asymptotic behavior of the probability distribution associated with the particular case of complete bipartite graphs, we argue that for this class of graphs, the number of rounds required to achieve a fixed approximation ratio must grow exponentially with the number of vertices/edges, a severe limitation on the performance of the Grover mixer. More than that, the construction suggests that it is likely that the same happens with other classes of graphs and even with other combinatorial optimization problems.
@@ II.1 The Grover-based QAOA
for a threshold value $t$ that must be optimized. The minus sign on Eq. (2) is because in this work we consider QAOA for minimization problems. The expectation value of GM-QAOA and GM-Th-QAOA of depth $r$ is denoted here by $E_{r}(\bm{\beta},\bm{\gamma})$ and $E_{r}(t)$  [^1] Despite the dependency of the expectation value concerning $\bm{\beta}$ and $\bm{\gamma}$, since we get the optimal angles at the beginning of the Sec. III, we simplified the notation from the outset., respectively, while $E_{r}$ denotes the expectation value of a generic Grover-based QAOA.
@@ II.2 Random variables
In this work, we apply the Headley and Wilhelm [28] approach, introduced for GM-QAOA, to the GM-Th-QAOA. That approach consists of using random variables as a mathematical model, writing the analytical expression of the expectation value in terms of the probability distribution associated with the solution space. If the probability distribution of a combinatorial optimization problem is known or can even be approximated, we can analytically optimize the angles $\bm{\beta}$ and $\bm{\gamma}$. That way, the expectation value can be computed directly from sampling the quantum circuit, avoiding the costly outer loop optimization procedure. For particular combinatorial optimation problems, Headley and Wilhelm [28] analytically conclude that the probability distribution of the Number Partition Problem with i.i.d. choice of the numbers follows a chi-squared distribution. On the other hand, in the empirical sense, it was observed that the solution space of the Capacitated Vehicle Routing and Portfolio Optimization problems seems to be normally distributed [22, 34].
In particular, Headley and Wilhelm [28] use continuous random variables to model the solution space as an asymptotic approximation on the large size limit. We, in contrast, use the exact case of discrete random variables. Formally, let $X$ be the random variable of uniformly sampling an element on the set $S$ and calculating the cost function. The function
$f_{X}(x)=\frac{|\{k\in S:c(k)=x\}|}{M}$ (3)
is the probability mass function (pmf) of $X$, and the support $R_{X}$ of $X$ is a countable subset of real numbers. We denote the mean, standard deviation, minimum value, and maximum value of $X$ by $\mu=\operatorname{E}[X]$, $\sigma=\sqrt{\operatorname{E}[X-\mu]^{2}}$, $R_{X}^{min}$, and $R_{X}^{max}$, respectively, and assume $0<\sigma<\infty$ (ignoring the degenerate distribution). Provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$, the approximation ratio is defined by
$\lambda=\frac{E}{R_{X}^{min}},\ E\in\{E_{r}(\bm{\beta},\bm{\gamma}),E_{r}(t),E_{r}\},$ (4)
that is, $E$ is the expectation value of the considered QAOA variant in each context. That metric, extensively used in the context of combinatorial optimization, measures the proximity of the solution provided by the algorithm to the optimal solution. We also define the random variable $Y$ and the standard random variable $Z$ as
$Y=X-\mu,\ Z=\frac{X-\mu}{\sigma}.$ (5)
For GM-QAOA, the summations of the original expectation value expression are replaced by the characteristic function of $X$ and its derivative (see Eq. 34 for the resulting analytical expressions). In contrast, for GM-Th-QAOA, due to the binary nature of the function $q(k)$ on that variant, the main statistical quantity of the analysis of GM-Th-QAOA is the conditional expectation. The expectation of $X$ given $X\leq x$ and the expectation of $X$ given $X>x$ are
@@
respectively, where
$G_{X}(x)=\sum_{k\in R_{X}:k\leq x}kf_{X}(k),\ F_{X}(x)=\sum_{k\in R_{X}:k\leq x}f_{X}(k).$ (7)
The function $F_{X}(x)$ is the cumulative distribution function (cdf). Note that with these definitions, we can consider any $x\in\mathbb{R}$ as argument. We replace, on the original expectation value expressions of GM-Th-QAOA, the statistical quantities $F_{X}(x)$ and $G_{X}(x)$, where the argument $x$ is the threshold value $t$, getting an expression that is dependent on the ratio between the number of states below or equal to the threshold and the total number of states.
Furthermore, in part of the present work, $F_{X}(x)$ and $G_{X}(x)$ must be differentiable. However, $X$ was defined to be discrete. To bypass that obstacle, note that we can write $F_{X}(x)$ and $G_{X}(x)$ as $F_{X}(x)=\sum_{k\in R_{X}}f_{X}(k)\theta(x-k)$ and $G_{X}(x)=\sum_{k\in R_{X}}kf_{X}(k)\theta(x-k)$, where $\theta(x)$ is the Heaviside step function. The derivatives of $F_{X}(x)$ and $G_{X}(x)$ are then given by $f_{X}^{G}(x)$ and $xf_{X}^{G}(x)$ respectively, where $f_{X}^{G}(x)=\sum_{k\in R_{X}}f_{X}(k)\delta(x-k)$, with $\delta(x)$ as the Dirac delta function. The distribution $f_{X}^{G}(x)$ is an extension of the concept of probability density function for non-continuous random variables, usually called generalized probability density function [43].
In other situations, on the other hand, it is convenient for $F_{X}(x)$ to be continuous, which does not occur for generalized density functions. That way, we consider $X$ as a continuous random variable with an asymptotic approximation, in the same way as Headley and Wilhelm [28]. Consequently, all summations are replaced by integrals, and the probability mass function becomes a probability density function (pdf). However, beyond mere convenience, this assumption is justifiable since the main QAOA target is NP-Hard optimization problems with sufficient scalability to make it a good approximation.